OpenStudy (anonymous):

solve on the interval [0,2pi) 1+sin theta=1/2 A. pi/6, 5pi/6 B. 2pi/3, 4pi/3 C. 7pi/6, 11pi/6 D. pi/3, 5pi/3

OpenStudy (anonymous):

Would the answer be C?

OpenStudy (anonymous):

\[1+\sin \theta =\frac{ 1 }{ 2 }\\ \sin \theta=\frac{ 1 }{ 2 }-1=-\frac{ 1 }{ 2 }\\now,\ given\ \theta \in \left[ 0 , 2\pi\right]\\ \implies \theta=\frac{ 7\pi }{ 6 },\frac{ 11\pi }{ 6}\\ [as\ per \ options, \ which \ satisfy \ interval\ and \ given\ eq.]\]